Jonah has been looking for a house. He found one he liked for $300,000. Jonah can afford a loan payment of $1,200 a month. He would like a 30-year loan. The bank is offering mortgage loans at a nominal annual interest rate of 6%, compounded monthly. How much of a down payment will Jonah need to make if he wants to buy this house?
Answer #1

Number of months = 12 x 30 = 360

Monthly interest rate = 6%/12 = 0.5%

PW of monthly payments ($) = 1,200 x P/A(0.5%, 360) = 1,200 x 166.7916** = 200,149.92

Down payment ($) = Loan amount - PW of monthly payments = 300,000 - 200,149.92 = 99,850.08

**P/A(0.5%, 360) = [1 - (1.005)-360] / 0.005 = (1 - 0.1660) / 0.005 = 0.8340 / 0.005 = 166.7916